# Full Line Code Completion: Bringing AI to Desktop

### v2

@@ Abstract
I Introduction
@@
In the past years, several products for multi-token code completion have appeared on the market, including Copilot [3], TabNine [4], Codeium [5], Amazon CodeWhisperer [6], Sourcegraph Cody [7], aiXcoder [8], and others. Most of such tools are cloud-based, so they miss a wide niche of customers, which includes users beyond a firewall, as well as users with limited internet access. While TabNine and aiXcoder support fully local models and data querying, they lack integration features and do not implement code correctness checks. This still leaves an open market for code completion engines that are local-based and ensure a proper user experience in an IDE.
In recent years, machine learning techniques were studied and applied for the different aspects of the completion problem [9, 10, 11, 12], including the generation of the completion sequences itself. Since source code can be interpreted as sequential data, natural language processing (NLP) methods have found their application for the task [13]. Specifically, Transformer-based [14] neural networks significantly outperform other architectures in the task of sequential data generation [15, 16, 17, 18].
@@ III-D1 Token healing
To address this issue in FLCC, we find an offset in the line, such that tokenization of the code to the left of the offset resembles training tokenization more. This involves backtracking left from the caret position for as long as the slice from the offset to the caret position is a full substring of one of the vocabulary tokens. Then, during model execution, we can filter out hypotheses that do not match characters that were cut out.
@@ IV-A Offline Evaluation
For offline evaluation, we use several repositories that are put into the test segment of the collected dataset. In these repositories, we pick a list of files of interest, and in each file, we generate various caret positions, where code completion is automatically invoked and evaluated. The picked repositories, files, and caret positions are fixed for every evaluation launch, so it is possible to track changes and study the model’s performance from iteration to iteration. We use an automatic IDE launch and programmatically invoke code completion. This allows us to use the entire pipeline implemented in the FLCC plugin, including caching, filtering, and other features. Moreover, this unlocks the ability not only to compare different versions of the plugin with each other but also to compare FLCC with standard code completion.
@@
Perfect lines. Ratio of FLCC invocations that fully match the actual code to the right of the position where it was invoked.
With such a setup, we run primary experiments for every major change implemented in FLCC. For example, while implementing our strategy for collecting hypotheses described in Section III-D2, for Python, we observed that its usage increased the percentage of Perfect Lines from 13% to 21%, while Matched Ratio only dropped from 20.6% to 20.3%, which is quite negligible.
@@ IV-B Online Evaluation
Offline evaluation is a valuable and convenient tool, but it might be biased, so it is not enough to make reliable decisions about product changes. As described in Section II-A, at JetBrains, we use early access programs (EAPs) to collect anonymous logs for experimental features and evaluate their performance. For ML features, in particular, we pay attention to rigorous A/B testing during EAPs. Such an approach allows to compare user experience with only standard completion versus both standard and Full Line Code Completion enabled. Also, this provides a convenient way of comparing different implementations of FLCC itself. During such A/B testing in late 2023 and 2024, we collected a lot of data to compute different important metrics from literature [62, 63]. We share the overall comparison for multiple languages in Table I and describe the metrics below, highlighting results for Python.
@@
For Python, the metric bumped 1.3 times (from 17% to 22%) for users with FLCC enabled compared to users with standard code completion only. The result is statistically significant with a p-value 0.001.
@@
The acceptance rate for Full Line Code Completion for Python was about 38%. For standard completion, the value is about 27%, however, no conclusions can be made from this comparison due to significantly different lengths of suggestions. Our solution does not perform as well as GitHub Copilot, which shows an acceptance rate of about 40% for both single-line and multi-line suggestions combined. However, one should note that GitHub Copilot uses a model that is two magnitudes larger and is cloud-based.
@@
The average time to show for FLCC is 150 ms, which is perfect for the user’s perception according to our knowledge. The result includes cache initialization events, as well as cache reusing events. According to our observation, the cache is used in over 90% of code completion events, which makes a typical time to show even lower.
@@ V Discussion
Also, the scope of the implemented feature is limited to a single line of code. On the one hand, this constraints the possible impact of such a time-saving feature, because it cannot generate bigger code snippets. On the other hand, the given limitation makes the feature controllable, meaning the relevance of the suggested line can be easily assessed by a developer. Additionally, by providing such a feature as a built-in functionality, we might raise the further adoption of AI-driven tools in the community.
Challenges and future directions. Despite the evident advantages, implementing FLCC in a local environment presented unique challenges, especially considering the constraints of memory footprint and processing speed. While we overcame these challenges, they highlight the need for optimizing ML models for practical use in constrained environments.
Additionally, the feature’s performance is significantly different for different languages, so we are aiming to study these differences and gain more insights about code completion for all those languages.
Bridging the gap between research and practice. Our work with FLCC also provides valuable lessons about the integration of complex, research-based solutions into commercial products. Our experience can serve as a guide for both academia and industry, highlighting a gap that often exists between theoretical model performance and practical deployment. In this regard, we hope that the insights we highlighted can be of use to future researchers. Furthermore, the developed gray text API and our approach to evaluating FLCC based on users’ data could be a model for future endeavors in the field, ensuring that such advanced tools not only pass the technical benchmarks but also align with user needs.
@@ VI Threats to Validity
State of the art. In contrast to most approaches in academic research, our focus extends beyond merely training a neural network for the code completion task. We aimed to develop the code completion feature, taking into consideration the behavioral patterns and preferences of users interacting with an IDE and its built-in code completion functionality. It is important to note that our analysis is limited to a specific group of users — those participating in the EAP. This limitation restricts our ability to conduct a direct comparison with other multi-token completion tools, which are commonly available as IntelliJ plugins and are frequently used by a broader user base not involved in the EAP.

### v3

@@ Abstract
Index Terms:
code completion, LLMs, IDE, online evaluationI Introduction
@@
In the past years, several products for multi-token code completion have appeared on the market, including Copilot [3], TabNine [4], Codeium [5], Sourcegraph Cody [6], aiXcoder [7], and others. Most of such tools are cloud-based, so they miss a wide niche of customers, which includes users beyond a firewall, as well as users with limited internet access. While TabNine and aiXcoder support fully local models and data querying, they lack integration features and do not implement code correctness checks. This still leaves an open market for code completion engines that are local-based and ensure a proper user experience in an IDE.
In recent years, machine learning techniques were studied and applied for the different aspects of the completion problem [8, 9, 10, 11], including the generation of the completion sequences itself. Since source code can be viewed as sequential data, natural language processing (NLP) methods have found their application for the task [12]. Specifically, Transformer-based [13] neural networks significantly outperform other architectures in generating sequential data [14, 15, 16, 17].
@@ III-D1 Token healing
To address this issue in FLCC, we find an offset in the line, such that tokenization of the code to the left of the offset resembles training tokenization more. This involves backtracking left from the caret position for as long as the slice from the offset to the caret position is a full substring of one of the vocabulary tokens. Then, during model execution, we can filter out hypotheses that do not match the cut-out characters.
@@ IV-A Offline Evaluation
For offline evaluation, we use several repositories from the test segment of the collected dataset. In them, we pick a list of files of interest, and in each file, we generate various caret positions, where code completion is automatically invoked and evaluated. The picked repositories, files, and caret positions are fixed for every evaluation launch, so it is possible to track changes and study the model’s performance from iteration to iteration. We use an automatic IDE launch and programmatically invoke code completion. This enables us to use the entire pipeline implemented in the FLCC plugin, including caching, filtering, and other features. Moreover, this allows not only to compare different versions of the plugin, but also to compare FLCC with standard code completion.
@@
Perfect lines. Ratio of FLCC invocations that fully match the actual code to the right of the invocation position.
With this setup, we run experiments for every major change implemented in FLCC. For example, while implementing our strategy for collecting hypotheses described in Section III-D2, we observed that, for Python, it increased the percentage of Perfect Lines from 13% to 21%, while Matched Ratio only dropped from 20.6% to 20.3%, which is quite negligible.
@@ IV-B Online Evaluation
Offline evaluation is a valuable and convenient tool, but it might be biased, so it is not enough to make reliable decisions about product changes. As described in Section II-A, at JetBrains, we use early access programs (EAPs) to collect anonymous logs for experimental features and evaluate their performance. For ML features, in particular, we pay attention to rigorous A/B testing during EAPs. Such an approach allows to compare user experience with only standard completion versus both standard and Full Line Code Completion enabled. Also, this provides a convenient way of comparing different implementations of FLCC itself. During such A/B testing in late 2023 and 2024, we collected a lot of data to compute different important metrics from literature [62, 63, 64]. We share the overall comparison for multiple languages in Table I and describe the metrics below, highlighting results for Python. All the provided comparisons are statistically significant, measured by employing per-user bootstrap [65].
We would like to emphasize that the choice of the proper online metric is a difficult task. While choosing the metrics, we paid attention to two goals. Firstly, metrics should be easy to compute and well-adopted in the community to put FLCC on the scale with other available code completion solutions. To that end, we employed the acceptance rate (AR). Secondly, we wanted to incorporate metrics that are agnostic to the source of code completion or a particular UX implementation (e.g., lookup completion or gray text completion). Thus, we used the ratio of completed code (RoCC) to be able to compare IDE code completion with and without FLCC.
@@
For Python, the metric bumped 1.3 times (from 17% to 22%) for users with FLCC enabled compared to users with standard code completion only.
@@
The acceptance rate for Full Line Code Completion for Python was about 38%. For standard completion, the value is about 27%, however, no conclusions can be made from this comparison due to significantly different lengths of suggestions. For comparison, GitHub Copilot shows an acceptance rate of about 30% [66] for both single-line and multi-line suggestions combined. However, one should be careful when comparing LLM-based completions using AR only, because this metric does not account for suggestion length or performance.
@@
The average time to show for FLCC is 150 ms, which is perfect for the user’s visual perception [67]. The result includes cache initialization events, as well as cache reusing. According to our observation, the cache is used in over 90% of completion events, which makes a typical time to show even lower.
@@ V Discussion
Also, the scope of the implemented feature is limited to a single line of code. On the one hand, this constraints the possible impact of the feature, because it cannot generate bigger code snippets. On the other hand, the given limitation makes the feature controllable, meaning the relevance of the suggested line can be easily assessed by a developer. Additionally, by providing such a feature as a built-in functionality, we might raise the further adoption of AI-driven tools in the community.
Challenges and future directions. Despite the evident advantages, implementing FLCC in a local environment presented unique challenges, especially considering the constraints of memory footprint and processing speed. While we overcame these challenges, they highlight the need for optimizing ML models for practical use in constrained environments. Additionally, the feature’s performance is significantly different for different languages, so we are aiming to study these differences and gain more insights about them.
Bridging the gap between research and practice. Our work with FLCC also provides valuable lessons about the integration of complex, research-based solutions into commercial products. Our experience can serve as a guide for both academia and industry, highlighting a gap that often exists between theoretical model performance and practical deployment. Furthermore, the developed gray text API and our approach to evaluating FLCC based on users’ data could be a model for future endeavors in the field, ensuring that such advanced tools not only pass the technical benchmarks but also align with user needs.
@@ VI Threats to Validity
State of the art. In contrast to most approaches in academic research, our focus extends beyond merely training a neural network for the code completion task. We aimed to develop the code completion feature, taking into consideration the behavioral patterns and preferences of users interacting with an IDE and its built-in code completion functionality. It is important to note that our analysis is limited to a specific group of users — those participating in the EAP. This limitation restricts our ability to conduct a direct comparison with other multi-token completion tools, which are used by a broader user base not involved in the EAP.